# ModalTune: Fine-Tuning Slide-Level Foundation Models with Multi-Modal Information for Multi-task Learning in Digital Pathology

### v1

@@ Abstract
Prediction tasks in digital pathology are challenging due to the massive size of whole-slide images (WSIs) and the weak nature of training signals. Advances in computing, data availability, and self-supervised learning (SSL) have paved the way for slide-level foundation models (SLFMs) that can improve prediction tasks in low-data regimes. However, working with these models is challenging, with issues such as catastrophic forgetting during fine-tuning and under-utilization of shared information between tasks and modalities. To overcome these two challenges, we propose ModalTune, a novel fine-tuning framework which introduces the Modal Adapter to integrate new modalities without modifying SLFM weights. Additionally, we use large-language models (LLMs) to encode labels as text, capturing semantic relationships and enhancing generalization across multiple tasks and cancer types in a single training recipe. ModalTune achieves state-of-the-art (SOTA) results against both uni-modal and multi-modal models across four cancer types, jointly improving survival and cancer subtype prediction while remaining competitive in pan-cancer settings. Additionally, we show ModalTune is highly generalizable to two out-of-distribution (OOD) datasets. To our knowledge, this is the first unified fine-tuning framework for multi-modal, multi-task, and pan-cancer modeling in digital pathology.
@@ 1 Introduction
Prediction tasks in histopathology are challenging due to the massive size and complexity of WSIs, which can contain critical diagnostic information for tumor staging, subtyping, and prognosis assessment. These predictive signals are often at the slide-level, requiring weakly-supervised learning approaches. Multiple Instance Learning (MIL) has become a key method for WSI analysis, dividing WSIs into patches, extracting features using pre-trained models, and aggregating them into slide-level representations through various pooling methods [29, 59, 10, 79, 61, 54]. The success of MIL is driven by the development of patch-level foundation models (FMs) [14, 69, 12, 44] for feature extraction, which enhance performance even in slide-level tasks [8].
Recent advances in computational resources, data availability, and algorithm efficiency have led to the development of SLFMs, which involve training MIL aggregators (‘slide encoders’) using SSL techniques. These SSL techniques include masked image modeling [27, 72], contrastive learning with transcriptomics [30, 64], pathology reports [19, 1], and multiple stains [33], marking a shift from patch-level FMs to robust SLFMs. However, these SLFMs cannot be directly applied with lightweight linear probing in real-world settings due to significant variations in WSI features caused by differences in tissue preparation across clinical centers, OOD morphological characteristics, and the presence of complex or rare diseases not seen during pre-training. These challenges require additional fine-tuning of SLFMs [17, 2] for deployment on new datasets.
Current frameworks for deploying SLFMs are limited to fine-tuning slide encoders on a single task and cancer site taking only WSIs as input, as shown in Fig. 1. They fail to fully leverage rich datasets, which may include other modalities like clinical records, reports, and genomics, as well as extensive slide annotations like cancer subtypes, tumor staging, prognosis, and biomarker presence. Existing frameworks face three key limitations when applied in this setting: (1) Standard fine-tuning often leads to catastrophic forgetting, causing pre-trained slide encoders to lose generalizable insights. (2) Existing SLFMs use only WSIs as input during inference, overlooking other modalities. While these models may have incorporated other modalities for training, there are significant benefits to leveraging them during inference to capitalize on multi-modal insights [32, 11]. (3) Current frameworks often emphasize single-task learning, overlooking interconnected multi-task information that could enhance SLFMs’ fine-tuning performance and generalization. [15]. These challenges result in suboptimally fine-tuned slide encoders.
@@
Our overall method called ModalTune is a flexible fine-tuning framework designed to fully leverage existing SLFMs and potential of pathology datasets. Our main contributions are as follows:
Modal Adapters: We propose Modal Adapters to mitigate catastrophic forgetting, maintain generalizability, and continually inject multi-modal information into slide encoders. It can interface with any transformer-based slide encoder and dynamically scale to new modalities.
@@ 3.1 Slide Encoder
We build on the SLFM, Gigapath [72], due to its strong performance and pre-trained weights. Gigapath effectively models slide embeddings using all patches and scales efficiently for cases involving multiple slides, aligning well with traditional MIL frameworks. However, our framework remains adaptable to any transformer-based slide encoder.
@@ 3.2.1 Transcriptomics Encoder
Bulk transcriptomics data is increasingly available and provides complementary insights on cancer biology to WSIs [9, 11, 32], allowing it to effectively predict patient survival [21, 6]. Thus, we consider it our additional modality in this work. However, we emphasize that any number of new modalities—such as other “omics”, imaging modalities, clinical data, or priors—can be integrated by concatenating their extracted feature tokens.
The transcriptomics encoder compresses paired bulk transcriptomics data, , containing genes into a set of tokens. We begin by dividing the genes into gene pathways, which are encoded using a sparse multi-layer perceptron (S-MLP) [21] into initial pathway features, , as described in [32]. To merge intra- and inter-pathway features, we use the MLP-mixer [62], encoding pathway features to produce . Finally, a linear projection reduces the number of pathways to , enhancing downstream tractability. The transcriptomics encoder ultimately outputs a compressed set of pathway tokens, .
@@ 3.3.1 Task Prompt
We anticipate that the Modal Adapter will require distinct feature sets tailored to each downstream task. To facilitate feature interactions and attention operations, we introduce a learnable [task] prompt. This consists of a set of learnable vectors, , for each downstream task , capturing task-specific biases across tasks. This prompt is concatenated with modal tokens—derived from compressed pathways in our approach—to create the complete set of multi-modal features: .
@@ 3.4 Text Embedding
We encode tasks using text to unify tasks, datasets, and cancer sites into a single embedding space, eliminating the need for complex loss functions. We generate text embeddings by converting clinical tabular data into text (Fig. 2) and processing it with a pre-trained text encoder. Here, we use the text encoder from CONCH [44], pre-trained on histology captions. Since different tasks require unique feature interactions (Sec. 3.3.1), we create a ‘general’ task () text embedding with all clinically relevant information and ‘task-specific’ () embeddings tailored to each downstream task. The text for the -th task is encoded into a vector, , and passed through a randomly initialized fixed projection layer to produce the final text embedding, . The supplementary provides further details on text construction (Sec. 9) and analysis (Sec. 10).
@@ 3.5 Training Objective
We aim to replicate the text embedding distribution to capture both task-specific information and semantic relationships. To achieve this, we use Kullback–Leibler (KL) divergence, that has demonstrated advantages over other distillation losses [41]. ModalTune is optimized to align the multi-modal embedding for task with its corresponding text embedding . The alignment is achieved by minimizing the KL divergence loss, averaged across all tasks, as shown in Eq. 5, where denotes softmax function.
@@ 4.1 Datasets
We utilized four TCGA datasets [63]—BRCA, NSCLC, GBMLGG, and RCC—to train and evaluate ModalTune for cancer subtype and survival prediction, covering 4,099 slides and 3,418 patients. COADREAD and BLCA datasets were used as OOD data to test model generalizability. Data was split at patient-level into [68%, 12%, 20%] for training, validation, and testing, stratified by cancer subtypes (Tab. 1). Transcriptomics data were sourced from the Xena database [25] and preprocessed similar to SurvPath [32], resulting in 331 pathways with 4,987 genes. For subtyping, subtypes were grouped into broader categories based on OncoTree code definitions [38], with rare subtypes ( cases) assigned to a ‘Rare-set’ used only for training with text embeddings. Baseline models cannot use this class due to insufficient occurrences for effective label-based training. Exact groupings are detailed in the supplementary materials (Sec. 8). For each patient, we extracted patches from WSI foreground regions at . Since outputs are generated on a patient-wise basis, patches from multiple WSIs were concatenated into a single bag. Features of size were extracted from each patch using Gigapath’s DINOv2 [48] pre-trained feature extractor.
@@ 4.3 ModalTune Setup
We configure ModalTune for tasks: general, survival, and subtype prediction. Gene pathways are linearly projected to dimensions for tractability, and the final output feature dimension is . Training runs for epochs with a learning rate and batch size of . Additional details are in supplementary (Sec. 7).
After training ModalTune, we evaluate its feature quality for classification and survival prediction. Using the General () task prompt, we extract features from both training and test datasets, selecting the best training epoch based on validation classification performance. For final alignment, we use linear probing, fitting a linear classifier [50] for cancer subtypes and a Cox Proportional Hazards (CPH) model [16] for survival prediction. All ModalTune results are multitask, utilizing a single feature set for both tasks. ‘ModalTune’ experiments are trained per cancer site and applied to downstream tasks within the same site, while ‘ModalTune Pan-Cancer’ pools all cancer sites during training and fits separate classifiers or CPH models per site for prediction.
@@ 4.4 Generalization Study Setup
BRCA GBMLGG NSCLC RCC Overall Cancer Subtype Prediction WSI ABMIL [29] TransMIL [59] Gigapath (Tuned) [72] Genomics MLP SNN [36] S-MLP [21] Gene Mixer Multi-modal MCAT [11] SurvPath [32] ABMIL (cat) [29] ABMIL (KP) [29] TransMIL (cat) [59] TransMIL (KP) [59] Gigapath (cat) [72] Gigapath (KP) [72] ModalTune (Ours) ModalTune Pan-Cancer (Ours) Survival Prediction WSI ABMIL [29] TransMIL [59] Gigapath (Tuned) [72] Genomics MLP SNN [36] S-MLP [21] Gene Mixer Multi-modal MCAT [11] SurvPath [32] ABMIL (cat) [29] ABMIL (KP) [29] TransMIL (cat) [59] TransMIL (KP) [59] Gigapath (cat) [72] Gigapath (KP) [72] ModalTune (Ours) ModalTune Pan-Cancer (Ours)
@@ 4.5 Quantitative Analysis
To address class imbalances in subtype classification, we report balanced accuracy, while survival prediction is evaluated using C-indices. Tab. 2 presents subtype classification and survival prediction results, showing the mean and standard deviation across three random seed runs. We report ablation studies on key design choices, including the effects of the Modal Adapter, text embeddings, text encoders, task prompts and projectors, in the supplementary (Sec. 11).
Overall Results: On average, we found ModalTune surpassed all other baselines on subtype classification and survival prediction tasks on multiple cancer sites, with and overall improvement, respectively, over the best baseline. ModalTune also greatly improved over the standard image-only SLFM framework, Gigapath ( in subtype, in survival prediction). These large improvements can be partially attributed to the introduction of the additional bulk transcriptomics modality. However, a fully-tuned Gigapath model with a late fusion of transcriptomics features (Gigapath (cat) and (KP)) still performs worse on classification and worse on survival prediction compared to ModalTune, indicating the benefits of our multi-task tuning framework (details in supp. Sec. 11.2). We believe the Modal Adapter’s ability to inject key multi-modal features without overfitting or catastrophic forgetting also contributed to this performance increase.
@@
Generalization Study: As shown in Tab. 3, the fully tuned Gigapath (cat) models severely overfit on BRCA, resulting in poor performance on both subtype and survival prediction tasks. These models fail to extract meaningful features for subtyping, defaulting to predicting the most common class (balanced accuracy of 0.5). In contrast, ModalTune Pan-Cancer demonstrated superior generalizability, performing only 2.4% worse than a fully-supervised Gigapath (cat) network in subtype prediction and actually outperforming it by 1.2% in survival prediction. ModalTune’s ability to extract highly relevant features without exposure to OOD datasets during fine-tuning underscores its capacity to maintain the generalizability of SLFMs. This result is particularly promising for the downstream application of ModalTune to smaller unseen datasets.
COADREAD BLCA Cancer Subtype Prediction Gigapath Sup. (cat) Gigapath Cls. (cat) Gigapath Surv. (cat) ModalTune ModalTune Pan-Cancer Survival Prediction Gigapath Sup. (cat) Gigapath Cls. (cat) Gigapath Surv. (cat) ModalTune ModalTune Pan-Cancer
@@ 4.6 Qualitative Analysis
To assess pathway contributions and modality interactions, we analyze two breast cancer cases from high and low-risk categories, as shown in Fig. 3. See supplementary for more qualitative analysis of t-SNE and Kaplan-Meier (Sec. 12).
@@ 5 Conclusion
In this study, we introduced ModalTune, a flexible fine-tuning framework that maximizes dataset usage for SLFMs in histopathology. To our knowledge, ModalTune is the first multi-modal, multi-task and pan-cancer tuning framework. By leveraging inter-modal and inter-task interactions, ModalTune outperformed all single-task baselines while preserving generalizability and mitigating catastrophic forgetting. Its resource-efficient variant, ModalTune Pan-Cancer, improves generalization on small OOD datasets, outperforming other fine-tuned models by a large margin while remaining competitive with supervised baselines. This underscores ModalTune’s ability to enhance SLFMs’ capacity for handling smaller datasets with rare diseases.
@@ 10 Text Embedding Analysis
Ablation BRCA GBMLGG NSCLC RCC Overall Cancer Subtype Prediction Single Modal No Text Embedding Single Task Prompt ModalTune with MedLlama v3.1 No Projector Trainable Projector Model-side Projector ModalTune 0.954 Survival Prediction Single Modal No Text Embedding Single Task Prompt ModalTune with MedLlama v3.1 No Projector Trainable Projector Model-side Projector ModalTune 0.750

### v2

@@ Abstract
Prediction tasks in digital pathology are challenging due to the massive size of whole-slide images (WSIs) and the weak nature of training signals. Advances in computing, data availability, and self-supervised learning (SSL) have paved the way for slide-level foundation models (SLFMs) that can improve prediction tasks in low-data regimes. However, current methods under-utilize shared information between tasks and modalities. To overcome this challenge, we propose ModalTune, a novel fine-tuning framework which introduces the Modal Adapter to integrate new modalities without modifying SLFM weights. Additionally, we use large-language models (LLMs) to encode labels as text, capturing semantic relationships across multiple tasks and cancer types in a single training recipe. ModalTune achieves state-of-the-art (SOTA) results against both uni-modal and multi-modal models across four cancer types, jointly improving survival and cancer subtype prediction while remaining competitive in pan-cancer settings. Additionally, we show ModalTune is generalizable to two out-of-distribution (OOD) datasets. To our knowledge, this is the first unified fine-tuning framework for multi-modal, multi-task, and pan-cancer modeling in digital pathology. Our code is publicly available at https://github.com/martellab-sri/ModalTune.
@@ 1 Introduction
Prediction tasks in histopathology are challenging due to the massive size and complexity of WSIs, which can contain critical diagnostic information for tumor staging, subtyping, and prognosis assessment. These predictive signals are often at the slide-level, requiring weakly-supervised learning approaches. Multiple Instance Learning (MIL) has become a key method for WSI analysis, dividing WSIs into patches, extracting features using pre-trained models, and aggregating them into slide-level representations through various pooling methods [28, 59, 9, 79, 61, 54]. With the development of large patch-level foundation models (FMs) [13, 69, 11, 43], feature extraction has improved, thereby enhancing the performance of MIL in slide-level tasks [7].
Recent advances in computational resources, data availability, and algorithm efficiency have led to the development of SLFMs, which involve training MIL aggregators (‘slide encoders’) using SSL techniques. These SSL techniques include masked image modeling [25, 72], contrastive learning with transcriptomics [29, 64], pathology reports [17, 1], and multiple stains [32], marking a shift from patch-level FMs to robust SLFMs. Yet, many domain-specific variations may still not be captured during pre-training, such as differences in tissue preparation across clinical centers, OOD morphological characteristics, and the presence of complex or rare diseases. This makes fine-tuning of SLFMs a more effective strategy for deployment on new datasets.
Current frameworks for deploying SLFMs are limited to fine-tuning them on a single task and cancer site while taking only WSIs as input (Fig. 1). They fail to fully leverage rich datasets, which may include other modalities like clinical records, reports, and genomics, as well as extensive slide annotations like cancer subtypes, tumor staging, prognosis, and biomarker presence. Existing frameworks face two key limitations when applied in this setting: (1) Existing SLFMs use only WSIs as input during inference, overlooking other modalities. While these models may have incorporated other modalities for training, there are significant benefits to leveraging them during inference to capitalize on multi-modal insights [31, 10]. (2) Current frameworks often emphasize single-task learning, overlooking interconnected multi-task information that could enhance SLFMs’ fine-tuning performance and generalization [14]. These challenges result in suboptimally fine-tuned SLFMs.
@@
Our overall method called ModalTune is a fine-tuning framework designed to fully leverage the potential of SLFMs and pathology datasets. Our main contributions are:
Modal Adapters: We propose Modal Adapters to continually inject multi-modal information into slide encoders without altering their weights. It can interface with any transformer-based slide encoder and dynamically scale to new modalities.
@@ 3.1 Slide Encoder
We build on the SLFM Gigapath [72], chosen for its strong performance, computational efficiency, and publicly available pre-trained weights. Gigapath effectively models slide embeddings by leveraging all patches and scales well in scenarios involving multiple slides with a large number of patches, enabling analysis without excessive overhead. This balance of efficiency and coverage makes it a practical choice for our study, even though more recent SLFMs may achieve stronger benchmark performance. Importantly, our framework remains fully adaptable to any transformer-based SLFM. For experiments on ModalTune applied to recent SLFM like TITAN [17], please refer to Supp. Sec. 12.
@@ 3.2.1 Transcriptomics Encoder
Bulk transcriptomics data is increasingly available and provides complementary insights on cancer biology to WSIs [8, 10, 31], allowing it to effectively predict patient survival [19, 5]. Thus, we consider it our primary additional modality in this work. Other modalities, such as clinical data (see Supp. Sec. 13), additional “omics,” imaging techniques, or domain-specific priors, can in principle be integrated by concatenating their extracted feature tokens. However, further exploration of these extensions is left for future work.
The transcriptomics encoder compresses bulk transcriptomics data, , containing genes into a set of tokens. We begin by dividing the genes into gene pathways, which are encoded using a sparse multi-layer perceptron (S-MLP) [19] into pathway features, , as described in [31]. To merge intra- and inter-pathway features, we use the MLP-mixer [62], encoding pathway features to produce . Finally, a linear projection reduces the number of pathways to , enhancing downstream tractability. The encoder finally outputs a compressed set of pathway tokens, .
@@ 3.3.1 Task Prompt
The Modal Adapter requires distinct feature sets tailored to each downstream task. To facilitate feature interactions and attention operations, we introduce a learnable [task] prompt. This consists of a set of learnable vectors, , for each downstream task , capturing task-specific biases across tasks. This prompt is concatenated with modal tokens derived from compressed pathways in our approach, to create the complete set of multi-modal features: .
@@ 3.4 Text Embedding
We encode tasks using text to unify tasks, datasets, and cancer sites into a single embedding space, eliminating the need for complex loss functions. We generate text embeddings by converting clinical tabular data into text (Fig. 2) and processing it with a pre-trained text encoder. Here, we use the text encoder from CONCH [43], pre-trained on histology captions. Since different tasks require unique feature interactions (Sec. 3.3.1), we create a ‘general’ task () text embedding with all clinically relevant information and ‘task-specific’ () embeddings tailored to each downstream task. The text for the -th task is encoded into a vector, , and passed through a fixed, randomly-initialized Projector to produce the final text embedding, . The Supplementary provides further details on text construction (Sec. 9) and analysis (Sec. 10).
@@ 3.5 Training Objective
We aim to replicate the text embedding distribution to capture both task-specific information and semantic relationships. To achieve this, we use Kullback–Leibler (KL) divergence, following [40], who demonstrated its advantages over CLIP [53] and other distillation losses. ModalTune is optimized to align the multi-modal embedding for task with its corresponding text embedding . The alignment is achieved by minimizing the KL divergence loss, averaged across all tasks, as shown in Eq. 5, where denotes softmax function.
@@ 4.1 Datasets
We utilized four TCGA datasets [63]—BRCA, NSCLC, GBMLGG, and RCC—to train and evaluate ModalTune for cancer subtype and survival prediction, covering 4,099 slides and 3,418 patients. COADREAD and BLCA datasets were used as OOD data to test model generalizability. Data was split at patient-level into [68%, 12%, 20%] for training, validation, and testing, stratified by cancer subtypes (Tab. 1). Transcriptomics data were sourced from the Xena database [23] and preprocessed similar to SurvPath [31], resulting in 331 pathways with 4,987 genes. For subtyping, subtypes were grouped into broader categories based on OncoTree code definitions [37], with rare subtypes ( cases) assigned to a ‘Rare-set’ used only for training with text embeddings. Baseline models cannot use this class due to insufficient occurrences for effective label-based training. Exact groupings are detailed in the Supp. Sec. 8. For each patient, we extracted patches from WSI foreground regions (via thresholding in HSV colorspace) at . Since outputs are generated on a patient-wise basis, patches from multiple WSIs were concatenated into a single bag. Features of size were extracted from each patch using Gigapath’s DINOv2 [47] pre-trained feature extractor.
@@ 4.3 ModalTune Setup
We configure ModalTune for tasks: general, survival, and subtype prediction. Gene pathways are linearly projected to dimensions for tractability, and the final output feature dimension is . Training runs for epochs with a learning rate and batch size of . Additional details are in Supp. Sec. 7.
After training, we evaluate its feature quality for classification and survival prediction. Using the General () task prompt, we extract features from both training and test datasets, selecting the best training epoch based on validation classification performance. For final alignment, we use linear probing (LP), fitting a linear classifier [49] for cancer subtypes and a Cox Proportional Hazards (CPH) model [15] for survival prediction. All ModalTune results are multitask, utilizing a single feature set for both tasks. ‘ModalTune’ experiments are trained per cancer site and applied to downstream tasks within the same site, while ‘ModalTune Pan-Cancer’ pools all cancer sites during training and fits separate classifiers or CPH models per site for prediction.
@@ 4.4 Generalization Study Setup
BRCA GBMLGG NSCLC RCC Overall Cancer Subtype Prediction WSI Gigapath LP [72] ABMIL [28] TransMIL [59] Gigapath (Tuned) [72] Genomics MLP SNN [35] S-MLP [19] Gene Mixer Multi-modal MCAT [10] SurvPath [31] ABMIL (cat) [28] ABMIL (KP) [28] TransMIL (cat) [59] TransMIL (KP) [59] Gigapath (cat) [72] Gigapath (KP) [72] ModalTune (Ours) ModalTune Pan-Cancer (Ours) Survival Prediction WSI Gigapath LP [72] ABMIL [28] TransMIL [59] Gigapath (Tuned) [72] Genomics MLP SNN [35] S-MLP [19] Gene Mixer Multi-modal MCAT [10] SurvPath [31] ABMIL (cat) [28] ABMIL (KP) [28] TransMIL (cat) [59] TransMIL (KP) [59] Gigapath (cat) [72] Gigapath (KP) [72] ModalTune (Ours) ModalTune Pan-Cancer (Ours)
@@ 4.5 Quantitative Analysis
To address class imbalances in subtype classification, we report balanced accuracy, while survival prediction is evaluated using C-indices. Tab. 2 presents subtype classification and survival prediction results, showing the mean and standard deviation across three random seed runs. We report ablation studies on key design choices, including Modal Adapters, text embeddings, text encoders, task prompts, and Projectors in the Supp. Sec. 11.
Overall Results: On average, we found ModalTune surpassed all other baselines on subtype classification and survival prediction tasks on multiple cancer sites, with and overall improvement, respectively, over the best baseline. ModalTune also greatly improved over the standard image-only SLFM framework, Gigapath (Tuned) ( in subtype, in survival prediction). These large improvements can be partially attributed to the introduction of the additional bulk transcriptomics modality. However, a fully-tuned Gigapath model with a late fusion of transcriptomics features (Gigapath (cat) and (KP)) still performs worse on classification and worse on survival prediction compared to ModalTune, indicating the benefits of our overall framework. Though TITAN is a stronger standalone model, applying ModalTune with it yielded similar trends (supp. Sec. 12).
@@
Generalization Study: As shown in Tab. 3, the fully tuned Gigapath (cat) models severely overfit on BRCA, resulting in poor performance on both subtype and survival prediction tasks. These models fail to extract meaningful features for subtyping, defaulting to predicting the most common class (balanced accuracy of 0.5). We also observe that ModalTune consistently outperforms Gigapath LP across both tasks and cancer sites. Furthermore, ModalTune Pan-Cancer demonstrated superior generalizability, performing only 2.4% worse than a fully-supervised Gigapath (cat) network in subtype prediction and actually outperforming it by 1.2% in survival prediction. ModalTune’s ability to extract highly relevant features without exposure to OOD datasets during fine-tuning underscores its capacity to maintain the generalizability of SLFMs. This result is particularly promising for the downstream application of ModalTune to smaller unseen datasets.
COADREAD BLCA Cancer Subtype Prediction Gigapath LP Gigapath Sup. (cat) Gigapath Cls. (cat) Gigapath Surv. (cat) ModalTune ModalTune Pan-Cancer Survival Prediction Gigapath LP Gigapath Sup. (cat) Gigapath Cls. (cat) Gigapath Surv. (cat) ModalTune ModalTune Pan-Cancer
@@ 4.6 Qualitative Analysis
To assess pathway contributions and modality interactions, we analyze two breast cancer cases from high and low-risk categories, as shown in Fig. 3. See Supp. for more qualitative analysis of t-SNE and Kaplan-Meier (Sec. 14).
@@ 5 Conclusion
In this study, we introduced ModalTune, a flexible fine-tuning framework that maximizes dataset usage for SLFMs in histopathology. To our knowledge, ModalTune is the first multi-modal, multi-task and pan-cancer tuning framework. By leveraging inter-modal and inter-task interactions, ModalTune outperformed all single-task baselines while retaining generalizability. Its resource-efficient variant, ModalTune Pan-Cancer, further improves generalization on small OOD datasets, outperforming other fine-tuned models by a large margin while remaining competitive with supervised baselines. This underscores ModalTune’s ability to enhance SLFMs’ capacity for handling smaller datasets.
@@ 10 Text Embedding Analysis
Ablation BRCA GBMLGG NSCLC RCC Overall Cancer Subtype Prediction Single Modal No Text Embedding Single Task Prompt ABMIL (cat) w/ text emb. ModalTune w/ LoRA ModalTune w/ MedLlama v3.1 No Projector Trainable Projector Model-side Projector ModalTune 0.954 Survival Prediction Single Modal No Text Embedding Single Task Prompt ABMIL (cat) w/ text emb. ModalTune w/ LoRA ModalTune w/ MedLlama v3.1 No Projector Trainable Projector Model-side Projector ModalTune 0.750
TITAN exp. BRCA GBMLGG NSCLC RCC Overall Cancer Subtype Prediction TITAN LP [17] TITAN (Tuned) [17] TITAN (cat) [17] TITAN (KP) [17] ModalTune TITAN 0.942 Survival Prediction TITAN LP [17] TITAN (Tuned) [17] TITAN (cat) [17] TITAN (KP) [17] ModalTune TITAN 0.735
Multimodal exp. BRCA GBMLGG NSCLC RCC Overall Cancer Subtype Prediction ModalTune 0.954 ModalTune w/ Clinical Survival Prediction ModalTune ModalTune w/ Clinical 0.755
COADREAD BLCA Cancer Subtype Prediction TITAN LP TITAN Sup. (cat) TITAN Cls. (cat) TITAN Surv. (cat) ModalTune TITAN Survival Prediction TITAN LP TITAN Sup. (cat) TITAN Cls. (cat) TITAN Surv. (cat) ModalTune TITAN